Working at home: According to the U.S Census Bureau, 34% of men who worked at home were college graduates. In a sample of 500 women who worked at home, 170 were college graduates.

(a) Find a point estimate for the proportion of college graduates among women who work at home. Round the answer to at least three decimal places.
Answer #1

Solution :

Given that,

n = 500

x = 170

Point estimate = sample proportion = = x / n = 170/500=0.340

Point estimate = = 0.340